The maximum applied force that can be applied to an 8.0 kg steel slider before it starts moving along a horizontal steel rail is 31 N [forward]. What is the coefficient of static friction?


1
Expert's answer
2021-10-18T10:59:06-0400

Applied force = Frictional force

"F= \mu_s mg \\

\mu_s = \frac{F}{mg} \\

= \frac{31}{8.0 \times 9.8} \\

= 0.3954"